OpenStudy (anonymous):

Hi, I am having the hardest time trying to figure out how in the example "Convert 2300 mm3 to m3?" the final answer was 2.3 × 10-6 m3. Will someone please help me?

OpenStudy (anonymous):

Hi there! So here's what's going on, I think... Your conversion might be going something like this... \[2300mm ^{3}\times \frac{ 1m }{ 1000mm } = 2.3m ^{3}\] (Nope!) That conversion factor is your problem. Converting between cubic units doesn't work quite like it does normally. Here's what you have to do: \[2300mm ^{3} \times \frac{ 1m * 1m * 1m }{ 1000mm * 1000mm *1000mm }\] Simplifying, that gets you: \[2300mm ^{3} \times \frac{ 1m ^{3} }{ 1.0*10^{9} mm ^{3} } = 2.3*10^{-6} m ^{3}\] Basically all you need to do is adjust your conversion factor to match what you're converting. Hope this helps :)